# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 2.2 Neural Network Class
To approximate the nuisance function $g_{o}(t,\mathbf{X})$, we use the popular class of sparse deep ReLU neural networks (schmidthieber2020nonparametric, and reference therein). This choice is motivated by two considerations: (i) ReLU networks are universal approximators capable of representing complex, high-dimensional functions; and (ii) their structure lends itself to sharp theoretical control through entropy bounds and sparsity constraints. We define a general $(K+1)$-layer feedforward ReLU network with depth $K\in\mathbb{N}_{+}$ and width vector $\boldsymbol{p}^{\top}=(p_{0},\ldots,p_{K+1})$, where $p_{0}=r$ denotes the input dimension and $p_{K+1}=1$ the scalar output. Specifically, the network is defined by
@@
which ensures controlled Lipschitz behavior of the class, where for a matrix $\bf A$} and a vector $\bf a$} the entrywise sup–norms is adopted, $\|\mathbf{A}\|_{\infty}=\max_{i,j}|A_{ij}|$, $\|\mathbf{a}\|_{\infty}=\max_{j}|a_{j}|$,. Following schmidthieber2020nonparametric we further impose sparsity. For sparsity level $s\in\mathbb{N}_{+}$ and range bound $D>0$, the sparse deep ReLU class is define as
@@
where $\|\cdot\|_{0}$ counts nonzero entries. Sparsity serves to control network complexity and is essential for obtaining optimal convergence rates. In the subsequent analysis we set the depth, width, and sparsity levels to balance bias-variance tradeoff. The bracketing entropy (vdvwellner1996, Chs. 2–3) of this class is then bounded using results of schmidthieber2020nonparametric, and Dudley’s entropy integral (vdvwellner1996, Thm. 2.14.1) is used to establish the required uniform empirical-process bounds.
3 FLEXI-Haz Procedure
Estimation is based on maximizing the full likelihood induced by the hazard model (1). Under noninformative and conditionally independent censoring given $({\bf X},{\bf Z})$, the corresponding full log–likelihood is proportional to
$\ell_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\Big[\Delta_{i}\big\{g(T_{i},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\int_{0}^{\tau}Y_{i}(t)\,\exp\!\big\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}\,dt\Big].$ (2)
where $Y_{i}(t)\;=\;\mathbf{1}(T_{i}\geq t)$, $t\in[0,\tau]$, $i=1,\ldots,n$, are at-risk processes, and $\tau$ is a prespecified constant. Because $g$ is modeled by a DNN, the predictable cumulative hazard in (2) contains a time-dependent integral that lacks a closed-form expression and must therefore be approximated numerically during optimization. We therefore introduce a grid $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ and approximate the integral by the Riemann–type sum
@@ 3 FLEXI-Haz Procedure
$\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\sum_{j:\,Y_{i}(t_{j-1})=1}\Big[\Delta_{ij}\,\big\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\big\{\min(T_{i},t_{j})-t_{j-1}\big\}\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\Big].$ (3)
In practice, discretizing the time axis on a fixed fine grid and augmenting it with all observed times (event and censoring times) provides an accurate and computationally efficient approximation, as detailed in Section 5.1. This construction automatically produces a dense grid in regions where failures occur, precisely where the integrand contributes most to the likelihood.
@@
$(\widehat{}\mbox{${\theta}$},\widehat{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\ell_{n}(\mbox{${\theta}$},g)\,.$ (4)
@@
$(\widetilde{}\mbox{${\theta}$},\widetilde{g})\;=\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta}\;\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\,.$ (5)
@@
Theorems 1–4 establish the asymptotic properties of (4), while Theorem 5 shows that under the chosen grid these properties carry over to the estimator (5) implemented in practice.
4 Asymptotic Analysis
4.1 Notation
@@
Some additional notation will be used throughout. We write $a_{n}\lesssim b_{n}$ if there exists a constant $c>0$ such that $a_{n}\leq c\,b_{n}$ for all $n$, and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. For a column vector $\bf a$}, ${\bf a}^{\otimes 2}={\bf a}{\bf a}^{\sf T}$. Let $P_{0}$ denote the distribution of a generic observation $(T,\Delta,{\bf X},{\bf Z})$. For any measurable function $h=h(T,{\bf X},{\bf Z})$ we define the unweighted $L^{2}_{P_{0}}$ norm $\|h\|^{2}_{L^{2}({P_{0}})}=\mathbb{E}_{0}\{h(T,{\bf X},{\bf Z})^{2}\}$ and the event–weighted norm $\|h\|^{2}_{L^{2}(P_{0},\Delta)}=\mathbb{E}_{0}\{\Delta\,h(T,{\bf X},{\bf Z})^{2}\big\}$.
Let $\eta=(\mbox{${\theta}$},g)$, write
@@ 4.2 Assumptions
The failure time $U$ and censoring time $C$ are conditionally independent given $({\bf X}^{\top},{\bf Z}^{\top})$, and the censoring mechanism is noninformative in the sense that the censoring process is noninformative for $(\mbox{${\theta}$},g)$. The true nuisance function satisfies $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$.
@@
The covariate vector $({\bf X},\mathbf{Z})$ takes values in a bounded subset of $\mathbb{R}^{p+q}$, denoted $\mathcal{X}\times\mathcal{Z}$, where both sets are compact. Without loss of generality, assume $\mathcal{X}\subset[0,1]^{d}$ and that $\|\mathbf{Z}\|_{2}\leq B_{{\bf Z}}$ almost surely. The joint density of $({\bf X},{\bf Z})$ is bounded away from zero on $\mathcal{X}\times\mathcal{Z}$. The parameter space for the finite-dimensional component is restricted to the compact set $\Theta\;=\;\{\vartheta\in\mathbb{R}^{p}:\ \|\vartheta\|\leq M_{\Theta}\}$. Finally, the study follows participants over a fixed, finite time horizon $\tau>0$.
@@
The network architecture satisfies $K=O(\log n)$, $s=O\!\big(n\gamma_{n}^{2}\log n\big)$, and $n\gamma_{n}^{2}\;\min_{0\leq\ell\leq K}p_{\ell}\lesssim\max_{0\leq\ell\leq K}p_{\ell}\lesssim n$. In addition, the sieve is uniformly bounded, $\|g\|_{\infty}\leq D_{n}$ for a known sieve radius $D_{n}=O(1)$.
@@
There exist $\delta_{1},\delta_{2}\in(0,1]$ such that, almost surely, $\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\geq\delta_{1}$ and $\Pr(U\geq\tau\mid{\bf X},{\bf Z})\geq\delta_{2}$.
@@
The covariate vector $\mathbf{Z}$ does not include a constant (intercept) term. Furthermore, the residual covariance matrix of $\mathbf{Z}$ given $\mathbf{X}$ is positive definite, that is,
$\Sigma_{\mathbf{Z}}=\mathbb{E}\!\left[\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid\mathbf{X})\big\}^{\top}\right]\succ 0\,.$
Assumptions A1A3, A5 used also in zhong2022deep under the partial linear Cox regression model. A3 determines the structure of the neural network and provides a trade-off between the approximation error and estimation error (schmidthieber2020nonparametric). Assumption A2 is a standard assumption for semi- or non-parametric regression. Moreover, A1 and A2 imply that there exist constants $0<c<C<\infty$ (that depend on $D_{n},B_{{\bf Z}}$ and $M_{\Theta}$) such that for all $t\in(0,\tau]$ and $({\bf X},{\bf Z})\in\mathcal{X}\times\mathcal{Z}$, $c_{h}\leq h_{o}(t|{\bf X},{\bf Z})\leq C_{h}$. Assumption A5 is a standard idenfiability assumption in survival analysis. Assumption A4 is required since in the hazard function (1) time is one of the components in the function $g_{o}$ estimated by the DNN and therefore the effects of the covariates $\bf X$}} and time are not disatengled, in contrast to the partial linear Cox model of zhong2022deep. Hence we need a non-vanishing chance to observe events over the entire study horizon for each covariate pattern.

### v2

@@ 2.2 Neural Network Class
To approximate the nuisance function $g_{o}(t,\mathbf{X})$, we use the popular class of sparse deep ReLU neural networks (Schmidt-Hieber, 2020, and references therein). This choice is motivated by two considerations: (i) ReLU networks are universal approximators capable of representing complex, high-dimensional functions; and (ii) their structure lends itself to sharp theoretical control through entropy bounds and sparsity constraints. We define a general $(K+1)$-layer feedforward ReLU network with depth $K\in\mathbb{N}_{+}$ and width vector $\boldsymbol{p}^{\top}=(p_{0},\ldots,p_{K+1})$, where $p_{0}=r=d+1$ denotes the input dimension and $p_{K+1}=1$ the scalar output. Specifically, the network is defined by
@@
which ensures controlled Lipschitz behavior of the class, where for a matrix $\bf A$} and a vector $\bf a$} the entrywise sup–norms are adopted, $\|\mathbf{A}\|_{\infty}=\max_{i,j}|A_{ij}|$, $\|\mathbf{a}\|_{\infty}=\max_{j}|a_{j}|$. Following Schmidt-Hieber (2020) we further impose sparsity. For sparsity level $s\in\mathbb{N}_{+}$ and range bound $D>0$, the sparse deep ReLU class is defined as
@@
where $\|\cdot\|_{0}$ counts nonzero entries. Sparsity serves to control network complexity and is essential for obtaining optimal convergence rates. In the subsequent analysis we set the depth, width, and sparsity levels to balance the bias-variance tradeoff. The bracketing entropy (van der Vaart and Wellner, 1996, Chs. 2–3) of this class is then bounded using results of Schmidt-Hieber (2020), and Dudley’s entropy integral (van der Vaart and Wellner, 1996, Thm. 2.14.1) is used to establish the required uniform empirical-process bounds.
3 FLEXI–Haz Procedure
3.1 Estimation of $(\mbox{${\theta}$}_{o},g_{o})$
Estimation is based on maximizing the full likelihood induced by the hazard model (2.1). Under noninformative and conditionally independent censoring given $({\bf X},{\bf Z})$, the corresponding full log–likelihood is proportional to
$\ell_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\Big[\Delta_{i}\big\{g(T_{i},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\int_{0}^{\tau}Y_{i}(t)\,\exp\!\big\{g(t,\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}\,dt\Big].$ (3.1)
where $Y_{i}(t)\;=\;\mathbf{1}(T_{i}\geq t)$, $t\in[0,\tau]$, $i=1,\ldots,n$, are at-risk processes, and $\tau$ is a prespecified constant. Because $g$ is modeled by a DNN, the predictable cumulative hazard in (3.1) contains a time-dependent integral that lacks a closed-form expression and must therefore be approximated numerically during optimization. We therefore introduce a grid $0=t_{0}<t_{1}<\cdots<t_{m_{n}}=\tau$ and approximate the integral by the Riemann–type sum
@@ 3.1 Estimation of $(\mbox{${\theta}$}_{o},g_{o})$
$\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\;=\;\frac{1}{n}\sum_{i=1}^{n}\sum_{j:\,Y_{i}(t_{j-1})=1}\Big[\Delta_{ij}\,\big\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\big\}-\big\{\min(T_{i},t_{j})-t_{j-1}\big\}\,\exp\!\{g(t_{j},\mathbf{X}_{i})+\mbox{${\theta}$}^{\top}\mathbf{Z}_{i}\}\Big].$ (3.2)
In practice, discretizing the time axis on a fine grid and augmenting it with all observed times (event and censoring times) provides an accurate and computationally efficient approximation, as detailed in Section 5.1. This construction automatically produces a dense grid in regions where failures occur, precisely where the integrand contributes most to the likelihood.
@@
$(\widehat{}\mbox{${\theta}$},\widehat{g})\;\in\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\ell_{n}(\mbox{${\theta}$},g)\,.$ (3.3)
@@
$(\widetilde{}\mbox{${\theta}$},\widetilde{g})\;\in\;\arg\max_{\,g\in\mathcal{G}(K,s,\boldsymbol{p},D),\;\theta\in\Theta\subseteq\mathbb{R}^{p}}\;\widetilde{\ell}_{n}(\mbox{${\theta}$},g)\,.$ (3.4)
@@
If the maximizer is not unique, owing to the non-uniqueness of neural-network representations, we take an arbitrary element of the argmax set. Theorems 1–4 establish the asymptotic properties of (3.3), while Theorem 5 shows that under the chosen grid these properties carry over to the estimator (3.4) implemented in practice.
3.2 Estimation and Marginal Inference for the Survival of a New Subject
We now turn to estimation and inference on the survival function of a new subject. Let $({\bf X}_{0},{\bf Z}_{0})$ be an independent draw from the covariate distribution, independent of the training sample, and define
$H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})\,du\,,\qquad S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})\}\,.$
[…]
$\widehat{H}(t\mid{\bf X}_{0},{\bf Z}_{0})=\sum_{j:t_{j}\leq t}\Delta t_{j}\exp\!\bigl\{\widetilde{g}(t_{j},{\bf X}_{0})+\widetilde{\mbox{${\theta}$}}^{\sf T}{\bf Z}_{0}\bigr\}\,$,
where $\Delta t_{j}=t_{j}-t_{j-1}$, and $\widehat{S}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-\widehat{H}(t\mid{\bf X}_{0},{\bf Z}_{0})\}$. Although the plug-in estimator is natural, it is generally not the most convenient starting point for first-order inference. The reason is that the target depends on the infinite-dimensional nuisance function $g_{o}$, and the first-order effect of nuisance estimation error need not vanish automatically. In particular, the nuisance-rate theory established in Section 4 controls $\widetilde{g}-g_{o}$ in an integrated norm, which is sufficient for estimation of $\mbox{${\theta}$}_{o}$ but does not by itself guarantee that the plug-in survival estimator admits an asymptotically linear expansion with an $o_{p}(n^{-1/2})$ remainder.
The purpose of the one-step construction is to remove this first-order nuisance bias (Bickel et al., 1993; Van der Vaart, 2000). Concretely, we will augment the plug-in estimator by an empirical average of an estimated efficient influence function (EIF). This correction yields an estimator whose leading error term is linear and whose remainder is second-order under the rate conditions imposed below. As a consequence, Wald-type inference becomes available for the cumulative hazard and, via the delta method, for the survival function.
In practice, we implement this correction using cross-fitting (Zheng and van der Laan, 2011; Chernozhukov et al., 2018).. The motivation for cross-fitting is twofold. First, it separates nuisance estimation from influence-function evaluation, which substantially weakens empirical-process requirements. Second, it makes the one-step correction more robust to overfitting in flexible nuisance estimators such as DNN. For this reason, we formulate the estimator below in cross-fitted form from the outset.
Two additional points are worth emphasizing. First, the target considered here is marginal in the sense that it is evaluated at an independent draw $({\bf X}_{0},{\bf Z}_{0})$. This differs from conditional pointwise inference at a fixed covariate value $({\bf x}_{0},{\bf z}_{0})$, which may be nonregular without additional smoothness or localization assumptions. Second, the construction below also applies to the hazard model without a separate linear component, obtained by absorbing $\bf Z$} into the nuisance function $g_{o}$. In that case, the parametric score and information terms disappear, and the EIF reduces to its nuisance–Riesz component alone. Equivalently, one may replace $\bf X$} throughout by the enlarged covariate vector $({\bf X}^{\sf T},{\bf Z}^{\sf T})^{\sf T}$.
We first present the estimator at a high level, postponing all technical derivations to Section 5 below.
To define the one-step estimator, let $\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})$ denote the efficient influence function for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$, where ${\bf W}=(T,\Delta,{\bf X},{\bf Z})$, and let $\widehat{\phi}_{t}^{(-k)}({\bf W};{\bf X}_{0},{\bf Z}_{0})$ be its estimator computed using nuisance estimators trained without fold $k$. Let $\{\mathcal{I}_{k}\}_{k=1}^{K}$ be a partition of $\{1,\dots,n\}$ into folds. For $i\in\mathcal{I}_{k}$, the observation ${\bf W}_{i}$ is evaluated using nuisance estimators fitted on the complement of $\mathcal{I}_{k}$.
The cross-fitted one-step estimator of the cumulative hazard is
$\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\widehat{H}^{\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})+\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0}),$ (3.5)
where
$\widehat{H}^{\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\sum_{k=1}^{K}\frac{|\mathcal{I}_{k}|}{n}\widehat{H}^{(-k)}(t\mid{\bf X}_{0},{\bf Z}_{0})$
and the corresponding one-step estimator of the survival function is
$\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\bigl\{-\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\bigr\}.$ (3.6)
At an intuitive level, (3.5) augments the plug-in estimator by a correction term that removes the first-order effect of nuisance estimation. Under the conditions stated below, the resulting estimator is asymptotically linear. The practical construction therefore reduces to two ingredients: estimation of the nuisance objects entering the plug-in hazard; and estimation of the efficient influence function $\phi_{t}^{0}(\cdot;{\bf X}_{0},{\bf Z}_{0})$.
The one-step correction is not shape-constrained. Hence, in finite samples, $\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})$ need not be monotone nondecreasing in $t$, and $\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})=\exp\{-\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\}$ need not be monotone nonincreasing. This is not a problem for the pointwise asymptotic inference developed here, which is for each fixed $t$. If a monotone curve is desired for display, the estimated survival curve can be post-processed by monotone rearrangement or isotonic projection.
The next subsection describes the resulting estimator explicitly in the finite discrete-covariate case, where the influence function admits a closed-form representation. The explicit construction below focuses on the case where $bX$ has finite support, since in that setting the nuisance Riesz representer admits a closed form. A more general high-level formulation for arbitrary covariate distributions is possible, based on an abstract $L^{2}(P_{{\bf X},{\bf Z}})$-valued pathwise differentiability argument and an associated representer condition, but this leads to a less explicit estimator and is therefore deferred to the Appendix.
3.2.1 Explicit form under finite discrete covariates
We now specialize to the case where $\bf X$} has finite support $\mathcal{X}=\{{\bf x}^{(1)},\dots,{\bf x}^{(J)}\}$ with ${\Pr}({\bf X}={\bf x}^{(j)})=\pi_{j}>0$, $j=1,\dots,J$. In this setting, the nuisance Riesz representer underlying the efficient influence function admits a closed form, and therefore the one-step estimator becomes fully explicit.
For $u\in[0,\tau]$ and $j=1,\dots,J$, define
$A_{j}(u)=\mathbb{E}_{0}\!\left\{Y(u)h_{o}(u\mid{\bf X},{\bf Z})\mid{\bf X}={\bf x}^{(j)}\right\}.$
Then the nuisance part of the efficient influence function is represented by a weight function of the form
$\psi_{t}^{0}(u,{\bf X};{\bf X}_{0},{\bf Z}_{0})=\mathbf{1}(u\leq t)\mathbf{1}({\bf X}={\bf X}_{0})\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)}.$ (3.7)
Accordingly, the efficient influence function takes the form
$\phi_{t}^{0}({\bf W};{\bf X}_{0},{\bf Z}_{0})=\int_{0}^{t}\mathbf{1}({\bf X}={\bf X}_{0})\frac{h_{o}(u\mid{\bf X}_{0},{\bf Z}_{0})}{\Pr({\bf X}={\bf X}_{0})\,A_{{\bf X}_{0}}(u)}\,dM(u)+c_{0,t}({\bf X}_{0},{\bf Z}_{0})^{\sf T}{\bf I}(\mbox{${\theta}$}_{0})^{-1}\ell_{\theta_{0}}^{\ast}({\bf W}),$ (3.8)
where $M(t)=N(t)-\int_{0}^{t}Y(u)h_{o}(u\mid{\bf X},{\bf Z})\,du$ is the counting-process martingale under $\mathbb{P}_{0}$ with $N(t)=\Delta\mathbf{1}(T\leq t)$. The quantities $c_{0,t}({\bf X}_{0},{\bf Z}_{0})$, $I(\mbox{${\theta}$}_{o})$, and $\ell^{*}_{\theta_{o}}$ are defined formally in Section 5; for now, it suffices to note that they represent the sensitivity term, efficient information, and efficient score, respectively. Replacing the unknown quantities in (3.8) by their estimators yields an explicit estimator $\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})$ and therefore an explicit cross-fitted one-step estimator through (3.5)–(3.6).
This discrete-support case is especially convenient in practice because the Riesz representer is explicit. Hence the one-step correction can be implemented directly, without introducing an additional approximation step for the representer.
Let $\widehat{\sigma}_{H,t}^{2}({\bf X}_{0},{\bf Z}_{0})$ denote a consistent estimator of the asymptotic variance of $\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})$, for example the empirical variance of the cross-fitted influence-function estimates is given by
$\widehat{\sigma}_{H,t}^{2}({\bf X}_{0},{\bf Z}_{0})=\frac{1}{n}\sum_{k=1}^{K}\sum_{i\in\mathcal{I}_{k}}\left\{\widehat{\phi}_{t}^{(-k)}({\bf W}_{i};{\bf X}_{0},{\bf Z}_{0})\right\}^{2}.$ (3.9)
An asymptotic $(1-\alpha)$ confidence interval for $H_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is therefore
$\widehat{H}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{\sigma}_{H,t}({\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}},$ (3.10)
and, by the delta method, an asymptotic $(1-\alpha)$ confidence interval for $S_{o}(t\mid{\bf X}_{0},{\bf Z}_{0})$ is
$\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\pm z_{1-\alpha/2}\frac{\widehat{S}^{1,\mathrm{cf}}(t\mid{\bf X}_{0},{\bf Z}_{0})\widehat{\sigma}_{H,t}({\bf X}_{0},{\bf Z}_{0})}{\sqrt{n}}.$ (3.11)
The asymptotic validity of these intervals is established in Section 5.
[…]
4.1 Assumptions
@@
Some additional notation will be used throughout. We write $a_{n}\lesssim b_{n}$ if there exists a constant $c>0$ such that $a_{n}\leq c\,b_{n}$ for all $n$, and $a_{n}\asymp b_{n}$ if both $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$ hold. For a column vector $\bf a$}, ${\bf a}^{\otimes 2}={\bf a}{\bf a}^{\sf T}$. Let $\mathbb{P}_{0}$ denote the true distribution of a generic observation $(T,\Delta,{\bf X},{\bf Z})$. For any measurable function $\mu=\mu(T,{\bf X},{\bf Z})$, define the unweighted $L^{2}_{\mathbb{P}_{0}}$ norm $\|\mu\|^{2}_{L^{2}({\mathbb{P}_{0}})}=\mathbb{E}_{0}\{\mu(T,{\bf X},{\bf Z})^{2}\}$ and the event–weighted norm $\|\mu\|^{2}_{L^{2}(\mathbb{P}_{0},\Delta)}=\mathbb{E}_{0}\{\Delta\,\mu(T,{\bf X},{\bf Z})^{2}\big\}$. Let $\eta=(\mbox{${\theta}$},g)$, write
@@
The failure time $U$ and censoring time $C$ are conditionally independent given $({\bf X}^{\top},{\bf Z}^{\top})$, and the censoring mechanism is noninformative in the sense that the censoring process is noninformative for $(\mbox{${\theta}$},g)$. The true nuisance function satisfies $g_{o}\in\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$.
@@
The covariate vector $({\bf X},\mathbf{Z})$ takes values in a bounded subset of $\mathbb{R}^{d+p}$, denoted $\mathcal{X}\times\mathcal{Z}$, where both sets are compact. Without loss of generality, assume $\mathcal{X}\subset[0,1]^{d}$ and that $\|\mathbf{Z}\|_{2}\leq B_{{\bf Z}}$ almost surely. The joint density of $({\bf X},{\bf Z})$ is bounded away from zero on $\mathcal{X}\times\mathcal{Z}$. The parameter space for the finite-dimensional component is restricted to the compact set $\Theta\;=\;\{\vartheta\in\mathbb{R}^{p}:\ \|\vartheta\|\leq M_{\Theta}\}$. Finally, the study follows participants over a fixed, finite time horizon $\tau>0$.
@@
The network architecture satisfies $K=O(\log n)$, $s=O\!\big(n\gamma_{n}^{2}\log n\big)$, and $n\gamma_{n}^{2}\lesssim\;\min_{0\leq\ell\leq K}p_{\ell}\lesssim\max_{0\leq\ell\leq K}p_{\ell}\lesssim n$. In addition, the sieve is uniformly bounded, $\|g\|_{\infty}\leq D_{n}$ for a known sieve radius $D_{n}=O(1)$.
@@
There exist $\delta_{1},\delta_{2},\delta_{3}\in(0,1]$ such that, almost surely, $\Pr(\Delta=1\mid T,{\bf X},{\bf Z})\geq\delta_{1}$, $\Pr(U\geq\tau\mid{\bf X},{\bf Z})\geq\delta_{2}$, and $\Pr(Y(t)=1\mid{\bf X},{\bf Z})\geq\delta_{3}$ for all $t\in[0,\tau]$.
@@
The covariate vector $\mathbf{Z}$ does not include a constant (intercept) term. Furthermore, the residual covariance matrix of $\mathbf{Z}$ given $\mathbf{T,X}$ among failures, is positive definite, that is,
$\Sigma_{\mathbf{Z}}=\mathbb{E}\!\left[\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid T,\mathbf{X},\Delta=1)\big\}\big\{\mathbf{Z}-\mathbb{E}(\mathbf{Z}\mid T,\mathbf{X},\Delta=1)\big\}^{\top}\right]\succ 0\,.$
Assumptions A1A3 are also used in Zhong et al. (2022) under the partial linear PH regression model. A3 determines the structure of the neural network and provides a trade-off between the approximation error and estimation error (Schmidt-Hieber, 2020). Assumption A2 is a standard assumption for semi- or non-parametric regression. Moreover, A1 and A2 imply that there exist constants $0<c<C<\infty$ (that depend on $D_{n},B_{{\bf Z}}$ and $M_{\Theta}$) such that for all $t\in(0,\tau]$ and $({\bf X},{\bf Z})\in\mathcal{X}\times\mathcal{Z}$, $c_{h}\leq h_{o}(t|{\bf X},{\bf Z})\leq C_{h}$. Assumption A4 is required since in the hazard function (2.1) time is one of the components in the function $g_{o}$ estimated by the DNN and therefore the effects of the covariates $\bf X$}} and time are not disentangled, in contrast to the partial linear Cox model of Zhong et al. (2022). Hence we need a non-vanishing chance to observe events over the entire study horizon for all $({\bf x},{\bf Z})$ in the support. While this assumption is stronger than in standard Cox regression, it is natural in baseline-free hazard models and is comparable to positivity conditions commonly imposed in semiparametric and causal models with continuous time. Assumption A5 is the non-PH analogue of the standard partial-linear identifiability condition. Because the nuisance component is allowed to depend on $(t,{\bf X})$, the relevant residual variation for the linear covariates must be measured after conditioning on $(T,{\bf X})$ among failures, rather than on $\bf X$} alone.